give number of moles and number of particles for atoms in 15.0 g of neon, Ne (20.18 g/mol)
1
Expert's answer
2020-03-25T12:48:31-0400

Solution:

1) Calculate the number of moles:

n(Ne) = m(Ne) / M(Ne) = (15.0 g) / (20.18 g/mol) = 0.743 mol.

n(Ne) = 0.743 mol.


2) The number of particles of neon (Ne) can be calculated using the Avogadro constant (Na) and the amount of Ne in mole:

number of particles = Avogadro constant × amount (mol)

N(Ne) = Na * n(Ne);

N(Ne) = (6.022*1023 particles/mol) * (0.743 mol) = 4.474*1023 particles of neon.

N(Ne) = 4.474*1023 atoms of Ne.


Answer: 0.743 mol and 4.474*1023 atoms in 15.0 g of neon (Ne).